OpenStudy (anonymous):

Write the domain and the range of a function in interval notation. Domain? And range of function?

OpenStudy (anonymous):

\[f(x)=\sqrt{x-1}\]

OpenStudy (anonymous):

domain is \[x \ge 1\]

OpenStudy (anonymous):

because negative square roots are not there!!!

OpenStudy (anonymous):

my options for an answer are

OpenStudy (anonymous):

\[\infty,0 \]

OpenStudy (anonymous):

i'm sorry other way around

OpenStudy (anonymous):

or like\[-1 \infty\]